Willis & Company has 20 million shares of $1 par value common stock outstanding. The company believes that its current market price of $100 per share is too high and decides to execute a 4-for-1 forward stock split to lower the price.

How many shares will be outstanding following the stock split?
What will be the new par value per share?
Answer #1

Outstanding shares = 20 million * 4/1

= 80 million

New par value per share = $1,* 1/4

= 0.25 per share

Comment if you face any issues
Answer #2

a) Outstanding share after stock split = 20*4 = 80 million

b) New par value per share = 1/4 = $0.25 per share